How many moles of Ag are in 32.0 grams of Ag2O? mol
Answer #1


Molar mass of Ag2O,
MM = 2*MM(Ag) + 1*MM(O)
= 2*107.9 + 1*16.0
= 231.8 g/mol


mass(Ag2O)= 32.0 g

use:
number of mol of Ag2O,
n = mass of Ag2O/molar mass of Ag2O
=(32 g)/(2.318*10^2 g/mol)
= 0.1381 mol
This is number of moles of Ag2O

one mole of Ag2O contains 2 mole of Ag
use:
number of moles of Ag = 2 * number of moles of Ag2O
= 2 * 0.1381
= 0.2761
Answer: 0.276 mol